An electronic device is used to measure the amount of charge that can flow into a memory storage area. How many electrons have been accumulated into the memory if the total charge is Q = -17.75 micro-Coulombs (μC) ? (The charge of one electron is e = -1.602×10-19 C)
Answer #1

q = ne

17.75* 10^-6 = n * 1.6* 10^-19

n = 1.108* 10^14 electrons

======

Comment before rate in case any doubt, will reply for sure.. goodluck